OpenStudy (anonymous):

What is the amplitude of the function F(x) = 3cos x? 3 1.5 6

OpenStudy (anonymous):

amplitude =3

OpenStudy (anonymous):

amplitude is always positive and it is in front of sin, cos or tan or their inverse

OpenStudy (anonymous):

The amplitude should be 3. The amplitude always appears as the coefficient :)

terenzreignz (terenzreignz):

Hi :) The amplitude is the highest value a trignometric function (either sin or cos) could attain Here's the general rule \[\huge y=a\sin(px+q)+b\](sin may be replaced by cos \[\huge y=\color{blue}{a}\sin(px+q)+b\]This part gives you the amplitude, \(\large |\color{blue}{a}|\) \[\huge y=a\sin(\color{red}{p}x+q)+b\]This part gives you the period \(\Large T=\frac{2\pi}{\color{red}p}\) \[\huge y=a\sin(px+\color{green}q)+b\]This part gives you the horizontal shift, or phase shift \(\Large \left| \frac{\color{green}q}{\color{red}{p}}\right|\) units to the right, if \(\Large \frac{\color{green}q}{\color{red}{p}}\) is negative, and to the left if \(\Large \frac{\color{green}q}{\color{red}{p}}\) is positive \[\huge y=a\sin(px+q)+\color{violet}b\]Finally, this part gives you the vertical shift, \(\large \color{violet}b\) units upward if b is positive and downward if b is negative

terenzreignz (terenzreignz):

@TeenaMathew tan does not have an amplitude.

OpenStudy (anonymous):

oh yeah you're right. thanks for reminding me

terenzreignz (terenzreignz):

^.^